# Paul Wild    Lutz Schröder    Karla Messing    Barbara König    Jonas Forster

### v1

@@ 5.1 Algorithmic considerations
$\delta^{\mathsf{HK}}(d)(A,B)=\max(\adjustlimits{\sup}_{\mu\in A}{\inf}_{\nu\in B}\delta^{\mathsf{K}}(d)(\mu,\nu),\adjustlimits{\sup}_{\nu\in B}{\inf}_{\mu\in A}\delta^{\mathsf{K}}(d)(\mu,\nu)).$
As the map $\mu\mapsto\inf_{\nu\in B}\delta^{\mathsf{K}}(d)(\mu,\nu)$ is convex, the left supremum can instead be taken over $A_{0}$ without changing the value, and similarly we may take the right supremum over $B_{0}$. It is however not in general true that the infima may be taken over $B_{0}$ and $A_{0}$, respectively.
@@
Let $X=\{x,y,z\}$, and assume $d(x,y)=d(x,z)=d(y,z)=1$. Let $A_{0}=\{\mu_{0},\mu_{1}\}$, $B_{0}=\{\mu_{2},\mu_{3}\}$, $\mu_{0}=\frac{1}{3}\cdot x+\frac{1}{3}\cdot y+\frac{1}{3}\cdot z$, $\mu_{1}=\frac{2}{3}\cdot x+\frac{1}{3}\cdot y$, $\mu_{2}=\frac{2}{3}\cdot y+\frac{1}{3}\cdot z$, $\mu_{3}=\frac{2}{3}\cdot z+\frac{1}{3}\cdot x$. Then the minimal distance from $\mu_{1}$ to $B$ is witnessed by $\mu^{*}=\frac{1}{2}\cdot\mu_{2}+\frac{1}{2}\cdot\mu_{3}=\frac{1}{6}\cdot x+\frac{1}{3}\cdot y+\frac{1}{2}\cdot z$.
Figure 1: Illustration of Example˜5. The thick line segments correspond to the sets $A=\mathsf{conv}(A_{0})$ and $B=\mathsf{conv}(B_{0})$, while the shaded region shows the radius-$\frac{1}{2}$ ball around $\mu_{1}$, which is part of a regular hexagon centered at $\mu_{1}$.
@@ 5.2 Borel Measures
As the categorical Kantorovich and Wasserstein constructions are typically considered in the shape of liftings or lax extensions of set endofunctors, their probabilistic instances are restricted to dealing with discrete probability distributions by necessity. The (probabilistic) Kantorovich-Rubinstein duality, meanwhile, is known to hold for much larger classes of probability distributions, such as Borel measures on metric spaces [30]. In this section we show that this is also true for the convex powerset duality.
For a pseudometric space $(X,d)$ we denote by $\mathsf{Bor}(X,d)$ the set of Borel probability measures, i.e. the probability measures defined on the $\sigma$-algebra generated by the open balls $B^{d}_{\varepsilon}({x})$. Every convex combination of Borel measures is itself a Borel measure. We can therefore also define a functor $\mathcal{C}^{\mathsf{Bor}}\colon\mathbf{PMet}\to\mathbf{PMet}$ where $\mathcal{C}^{\mathsf{Bor}}(X,d)$ is the set of non-empty convex subsets of $\mathsf{Bor}(X,d)$, equipped with Hausdorff-Kantorovich distance $\delta^{\mathsf{HK}}(d)=\delta^{\mathsf{H}}(\delta^{\mathsf{K}}(d))$, where both $\delta^{\mathsf{H}}$ and $\delta^{\mathsf{K}}$ are defined as before. Note that every nonexpansive map on $(X,d)$ is Borel-measurable, so that no issues arise when taking expected values.
$\delta^{\mathsf{K}}(d)(\mu,\nu)=\sup\{|{\textstyle\int_{X}{f}\,\mathrm{d}{\nu}}-{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}|\mid f\colon(X,d)\to([0,1],d_{e})\text{ nonexp.}\}$
@@
Due to a general result by Goncharov et al. [13], we know that $\mathcal{C}^{\mathsf{Bor}}$ admits a Kantorovich representation for some suitable class of predicate liftings. By [13, Theorem 5.3] it suffices to show that it preserves initial morphisms. In the real-valued setting these correspond to isometries, i.e. nonexpansive maps $f\colon(X,d_{1})\to(Y,d_{2})$ such that $d_{2}(f(x),f(x^{\prime}))=d_{1}(x,x^{\prime})$ for all $x,x^{\prime}\in X$.
@@
The class of predicate liftings one obtains is quite large; we show that one can in fact make do with just a single predicate lifting, which strengthens the corresponding instance of the coalgebraic quantitative Hennessy-Milner theorem [13, Corollary 5.10] by providing a compact explicit syntax for the expressive logic:
@@
In the proof, we follow the same steps as before, first passing from pseudometrics to metrics and then leveraging linear algebra to obtain duality.
@@ 6 Conclusions and Future Work
We have proved generalized Kantorovich-Rubinstein duality, i.e. coincidence of coupling-based (i.e. Wasserstein) and price-function-based (i.e. Kantorovich or codensity) presentations of functor liftings induced by a given choice of modalities, for two important and non-trivial cases: the Lévy-Prokhorov distance on distributions, and the standard distance on convex sets of distributions that arises from composing the Hausdorff and Kantorovich-Wasserstein metrics. In both cases, we obtain a characterization of the respective distance by means of quantitative modal logics defined by the given modalities; for the case of the Lévy-Prokhorov distances, this logic is (up to restriction of the propositional base) the logic of generally previously studied in context of fuzzy description logics [27], and in the second case the involved modality is just the composite of the usual fuzzy diamond and the expectation modality [32]. In the case of convex powerset, we demonstrate additionally that the Kantorovich presentation plays out algorithmic advantages in the actual computation of distances.
We leave several key open problems, among them on the one hand the extension of the duality result for the Lévy-Prokhorov metric from discrete to Borel probability distributions (while we already prove the duality result for convex powerset at this level of generality), and on the other hand the extension of the duality result for convex powerset to unrestricted fuzzy relations in place of pseudometrics (already established in our result on the Lévy-Prokhorov metric). The latter generalization will amount to a duality result for the known Wasserstein lax extension of convex powerset [32]. Also, we aim to capitalize on the present result in the design of algorithms that actually compute distinguishing formulae as witnesses of lower bounds on behavioural distance, complementing recent results on behavioural distance under the Kantorovich-Wasserstein distance of distributions [24, 28].

### v2

@@ 5.1 Algorithmic considerations
$\delta^{\mathsf{HK}}(d)(A,B)=\max(\adjustlimits{\sup}_{\mu\in A}{\inf}_{\nu\in B}\delta^{\mathsf{KR}}(d)(\mu,\nu),\adjustlimits{\sup}_{\nu\in B}{\inf}_{\mu\in A}\delta^{\mathsf{KR}}(d)(\mu,\nu)).$
As the map $\mu\mapsto\inf_{\nu\in B}\delta^{\mathsf{KR}}(d)(\mu,\nu)$ is convex, the left supremum can instead be taken over $A_{0}$ without changing the value, and similarly we may take the right supremum over $B_{0}$. It is however not in general true that the infima may be taken over $B_{0}$ and $A_{0}$, respectively.
@@
Let $X=\{x,y,z\}$, and assume $d(x,y)=d(x,z)=d(y,z)=1$. Let $A_{0}=\{\mu_{0},\mu_{1}\}$, $B_{0}=\{\mu_{2},\mu_{3}\}$, $\mu_{0}=\frac{1}{3}\cdot x+\frac{1}{3}\cdot y+\frac{1}{3}\cdot z$, $\mu_{1}=\frac{2}{3}\cdot x+\frac{1}{3}\cdot y$, $\mu_{2}=\frac{2}{3}\cdot y+\frac{1}{3}\cdot z$, $\mu_{3}=\frac{2}{3}\cdot z+\frac{1}{3}\cdot x$. Then the minimal distance from $\mu_{1}$ to $B$ is witnessed by $\mu^{*}=\frac{1}{2}\cdot\mu_{2}+\frac{1}{2}\cdot\mu_{3}=\frac{1}{6}\cdot x+\frac{1}{3}\cdot y+\frac{1}{2}\cdot z$. See Figure˜1 for an illustration.
Figure 1: Illustration of Example˜5. The thick line segments correspond to the sets $A=\mathsf{conv}(A_{0})$ and $B=\mathsf{conv}(B_{0})$, while the shaded region shows the radius-$\frac{1}{2}$-ball around $\mu_{1}$, which is part of a regular hexagon centered at $\mu_{1}$.
@@ 5.2 Borel Measures
As the categorical coupling-based and price-function-based constructions are typically considered in the shape of liftings or lax extensions of set endofunctors, their probabilistic instances are restricted to dealing with discrete probability distributions by necessity. The (probabilistic) Kantorovich-Rubinstein duality, meanwhile, is known to hold for much larger classes of probability distributions, such as Radon measures on metric spaces [37]. In this section we show that this is also true for the convex powerset duality.
For a pseudometric space $(X,d)$ we denote by $\mathsf{Bor}(X,d)$ the set of Borel probability measures, i.e. the probability measures defined on the $\sigma$-algebra generated by the open balls $B^{d}_{\varepsilon}({x})$. Every convex combination of Borel measures is itself a Borel measure. We can therefore also define a functor $\mathcal{C}^{\mathsf{Bor}}\colon\mathbf{PMet}\to\mathbf{PMet}$ where $\mathcal{C}^{\mathsf{Bor}}(X,d)$ is the set of non-empty convex subsets of $\mathsf{Bor}(X,d)$, equipped with Hausdorff-Kantorovich distance $\delta^{\mathsf{HK}}(d)=\delta^{\mathsf{H}}(\delta^{\mathsf{KR}}(d))$, where both $\delta^{\mathsf{H}}$ and $\delta^{\mathsf{KR}}$ are defined as before. Note that every nonexpansive map on $(X,d)$ is Borel-measurable, so that no issues arise when taking expected values.
$\delta^{\mathsf{KR}}(d)(\mu,\nu)=\sup\{|{\textstyle\int_{X}{f}\,\mathrm{d}{\nu}}-{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}|\mid f\colon(X,d)\to([0,1],d_{e})\text{ nonexp.}\}$
@@
Due to a general result by Goncharov et al. [14], we know that $\mathcal{C}^{\mathsf{Bor}}$ admits a Kantorovich representation for some suitable class of predicate liftings. By [14, Theorem 5.3] it suffices to show that it preserves initial morphisms. In the real-valued setting these correspond to isometries, i.e. nonexpansive maps $f\colon(X,d_{1})\to(Y,d_{2})$ such that $d_{2}(f(x),f(x^{\prime}))=d_{1}(x,x^{\prime})$ for all $x,x^{\prime}\in X$.
@@
The class of predicate liftings one obtains is quite large; we show that one can in fact make do with just a single predicate lifting, which strengthens the corresponding instance of the coalgebraic quantitative Hennessy-Milner theorem [14, Corollary 5.10] by providing a compact explicit syntax for the expressive logic:
@@
In the proof, we follow the same steps as before, first passing from pseudometrics to metrics and then leveraging linear algebra to obtain coincidence of the two distances.
Remark 4.
LABEL:thm:convex-powerset-duality-borel establishes that the composition of the Hausdorff and Kantorovich-Rubinstein distances can be expressed in price-function form with respect to the single modality $A\mapsto\sup_{\mu\in A}{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}$. To extend this into a full-blown duality result for continuous measures, stating that $\delta^{\mathsf{HK}}$ also coincides with the corresponding coupling-based construction, two additional steps are needed: First, we need to have duality for the two individual distances. In the case of the Hausdorff distance, no additional requirements are necessary, but in the case of Kantorovich-Rubinstein distance one needs to restrict from Borel measures over general pseudometric spaces to one of the settings in which duality is known to hold, such as Polish spaces or Radon measures [37]. Second, one needs to establish a compositionality result for the respective coupling-based constructions, generalizing the one from the discrete case [39, Example 6.8.3]. We leave the details of such a duality result for future work.
@@ 6 Conclusions and Future Work
We have proved generalized Kantorovich-Rubinstein duality, i.e. coincidence of coupling-based (or Wasserstein) and price-function-based (or Kantorovich or codensity) presentations of functor liftings induced by a given choice of modalities, for two important and non-trivial cases: the Lévy-Prokhorov distance on distributions, and the standard distance on convex sets of distributions that arises from composing the Hausdorff and Kantorovich-Rubinstein metrics. In both cases, we obtain a characterization of the respective distance by means of quantitative modal logics defined by the given modalities; for the case of the Lévy-Prokhorov distances, this logic is (up to restriction of the propositional base) the logic of generally previously studied in context of fuzzy description logics [34], and in the second case the involved modality is just the composite of the usual fuzzy diamond and the expectation modality [39]. In the case of convex powerset, we demonstrate additionally that the price-function-based presentation plays out algorithmic advantages in the actual computation of distances.
We leave several key open problems, among them on the one hand the extension of the duality result for the Lévy-Prokhorov metric from discrete to Borel probability distributions, and on the other hand the extension of the duality result for convex powerset to unrestricted fuzzy relations in place of pseudometrics (already established in our result on the Lévy-Prokhorov metric). The latter generalization will amount to a duality result for the known coupling-based lax extension of convex powerset [39]. Also, we aim to capitalize on the present result in the design of algorithms that actually compute distinguishing formulae as witnesses of lower bounds on behavioural distance, complementing recent results on behavioural distance under the Kantorovich-Rubinstein distance of distributions [31, 35].